OpenStudy (anonymous):

A sequence has its first term equal to 8 and each term of the sequence is obtained by adding 6 to the previous term. If f(n) represents the nth term of the sequence, which of the following recursive functions best defines this sequence? f(1) = 6 and f(n) = f(n - 1) + 8; n > 1 f(1) = 8 and f(n) = f(n - 1) + 6; n > 1 f(1) = 8 and f(n) = f(n - 1) + 6n; n > 1 f(1) = 6 and f(n) = f(n - 1) + 8n; n > 1

OpenStudy (anonymous):

f(n-1) represents the previous term Then add 8 to the previous term, so f(n-1) + 8= f(n)

OpenStudy (anonymous):

Wait sorry! I mean to say f(n-1)+6-f(n)! And f(1) = 8

OpenStudy (anonymous):

f(n-1) + 6 = f(n)

OpenStudy (lena772):

B

OpenStudy (anonymous):

Correct!